OpenStudy (anonymous):

solve the following differential equation subject to indicated condition: dy/dx = 1 / (sqrt(x+1)) ; y=18 & x=3

OpenStudy (anonymous):

y' = (x+1)^(-1/2) u can just do normally since derivative of bracket is 1 y =( x+1)^1/2/(1/2) +c = 2 (x+1)^(1/2) +c and then sub y and x to get c.